# ReverseNER: A Self-Generated Example-Driven Framework for Zero-Shot Named Entity Recognition with Large Language Models

### v3

@@ 5 Conclusion
In conclusion, ReverseNER provides a sophisticated methodology for addressing zero-shot NER by leveraging self-generated examples within an example-driven framework. By constructing a labeled example library and employing self-consistency scoring, ReverseNER effectively mitigates the challenges associated with NER in the absence of annotated datasets. Experimental evaluations indicate substantial improvements over conventional zero-shot and few-shot NER approaches, underscoring its potential for enhancing entity recognition capabilities. Future research endeavors will focus on optimizing computational efficiency and broadening the applicability of ReverseNER to encompass more diverse NLP tasks.
6 Limitations
@@
  • The self-consistency scoring mechanism increases the volume of reasoning attempts, which may escalate runtime and incur additional costs, raising scalability concerns in resource-limited settings.
@@ 7 Acknowledgment
Appendix A Actual Prompts for Constructing the Example Library
Taking the WikiGold dataset as an example, the prompt used for generating entity vocabularies of this dataset is illustrated in Listing A1. We repeat this prompt when constructing example labels for each cluster.
The prompt for constructing example labels in the example library is shown in Listing A2. The entity vocabularies it uses are derived from the results of the corresponding entity vocabulary generation prompt. The reference sentence provided at the end corresponds to the sentence located at the center of the respective cluster. We repeat this prompt when constructing example labels for each cluster.
@@ Appendix A Actual Prompts for Constructing the Example Library
Listing 1: Vocabulary Generation Prompt
@@
Listing 2: Label Generation Prompt
Appendix B Actual Prompts for Performing NER on Each Task Input
The prompt used for inferring a specific input sentence from the WikiGold dataset is shown in Listing B3. The provided examples are selected from the constructed example library based on cosine similarity.

### v4

@@ 5 Conclusion
This paper presents ReverseNER, a novel method designed to enhance zero-shot named entity recognition (NER) tasks by utilizing the generative capabilities of large language models (LLMs) in a self-generated, example-driven approach.
Our experiments across multiple benchmarks demonstrate that ReverseNER consistently improves NER performance compared to other zero-shot methods with significantly lower resource consumption. Additionally, the incorporation of an entity-level self-consistency scoring mechanism further refines the precision of entity recognition, particularly in scenarios involving ambiguous entity types. The analysis highlights the effectiveness of our feature sentence extraction and example library construction processes, ensuring high-quality and diverse examples that significantly aid the LLM’s performance. This paper also presents three metrics to evaluate the quality of the generated example library to increase interpretability.
Overall, ReverseNER contributes to the ongoing efforts to leverage LLMs for more effective and autonomous information extraction, providing a practical approach for improving zero-shot NER performance in various applications.
6 Limitations and Future Work
@@
  • The effectiveness of the proposed method is highly dependent on the quality of manually provided entity type definitions.
@@
  • The optional self-consistency scoring mechanism increases the frequency of LLM invocations, which undermines the low-resource consumption advantage achieved through the reversed NER example library construction process.
This work primarily focuses on enhancing the performance of LLMs for NER tasks. However, the concept of guiding LLMs to self-hint through reversed processes holds the potential for improving performance in other NLP tasks, such as relation extraction, sentiment analysis, and semantic role labeling.
@@ 7 Acknowledgment
Appendix A Dataset Statistics
Dataset statistics are shown in Table 9. For WikiGold and GovAff, we perform the zero-shot NER task on the entire dataset. For CoNLL03 and People’s Daily, we restrict the zero-shot NER task to the official test sets. Since this work focuses on zero-shot NER, no training or validation data are utilized.
Table 9: Dataset Statistics
Dataset Sentences Tested in this Paper Full Dataset Size
CoNLL03 3453 14382
WikiGold 1697 1697
People’s Daily 3000 19484
GovAff 4825 4825
Appendix B Actual Prompts for Constructing the Example Library
Taking the WikiGold dataset as an example, the prompt used for generating entity vocabularies of this dataset is illustrated in Listing B1. We repeat this prompt when constructing entity-labeled examples for each cluster.
The prompt for constructing entity-labeled examples in the example library is shown in Listing B2. The entity vocabularies it uses are derived from the results of the corresponding entity vocabulary generation prompt. The reference sentence provided at the end corresponds to the sentence located at the center of the respective cluster. We repeat this prompt when constructing entity-labeled examples for each cluster.
@@ Appendix B Actual Prompts for Constructing the Example Library
Listing 1: Entity Vocabulary Generation Prompt
@@
Listing 2: Entity-labeled Sentences Construction Prompt
Appendix C Actual Prompts for Performing NER on Each Task Input
The prompt used for inferring a specific input sentence from the WikiGold dataset is shown in Listing C3. The provided examples are selected from the constructed example library based on cosine similarity.